At an angle of elevation 610 captain on a ship noticed a person on the top of a cliff waiving towards the ship. To see him better, captain moved his ship 92m closer to the cliff. The angle of elevation at that moment was 690 . Determine the height of the cliff to the nearest tenth of a meter.


1
Expert's answer
2020-11-08T18:05:21-0500

SOLUTION

Here is the sketch of the scenario;


First, we separate the triangles;



Here, we apply the Sine Rule;

"\frac{a}{sinA}=\frac{b}{sinB}=\frac{c}{sinC}"


In this case we apply a and c. i.e.


"\frac{a}{sinA}=\frac{c}{sinC}"


"\frac{92}{sin(8°)}=\frac{c}{sin(61°)}"


"c = \frac{92×sin(61°)}{sin(8°)}"


"c = 578.16m"


Therefore, "X = 578.16m"


Then, considering the last triangle, which is right-angle triangle, we have;


We consider the SOHCAHTOA concept, where we apply the sine formula (SOH)

"sin(69°)=\frac{H}{578.16}"


"H= 578.16 \u00d7 sin (69\u00b0)"


"H = 539.76m" (To the nearest tenths)

"H = 539.8m"

Therefore, Height of the cliff is "539.8m"